Two point charges are seperated by a distance of 100cm. With a magnitude of -10-6 for both charges. Find the electric force between them


1
Expert's answer
2022-04-17T17:06:55-0400

According to the Coulomb's law the force is:


"F = k\dfrac{|q_1q_2|}{r^2 }"

where "k=9\times 10^9 N\cdot m^2/C^2" is the Coulomb's constant, "q_1=q_2=-10^{-6}C" are the magnitude of the charges, and "r= 100cm = 1m" is the distance between them. Thus, obtain:


"F =9\times 10^9\cdot \dfrac{(-10^{-6})^2}{1^2 } = 9\times 10^{-3}N"

Answer. "9\times 10^{-3}N".